# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v2

@@ 9.1. The free energy of a toric log Fano curve
Denote by $\mathcal{H}_{1,1}(-K_{(X,\Delta)})^{S^{1}}$ the space of all $S^{1}$–invariant psh metrics $\phi$ on $K_{(X,\Delta)}$ such that $\phi(x)$ is in $C^{1,1}(\mathbb{R}).$ Given $\phi_{0}\in\text{PSH(}-K_{(X,\Delta})_{\mathrm{b}}$ which is assumed $S^{1}$–invariant, the corresponding free energy $F_{\gamma}(\mu)$ satisfies the following formula, when $\mu=dd^{c}\phi/V$ for $\phi\in\mathcal{H}(-K_{(X,\Delta)})_{1,1}^{S^{1}}$:
@@
When $\gamma=1$ it is shown in the proof of [16, Lemma 5, step 1] that
@@
(in our normalization $V\mathcal{M}$ corresponds to the functional $\mathcal{M}$ in [16]). By formula 3.13, this means that
@@
Note that, by formula 9.1, $D(\mu|\mu_{\phi})=\int_{\mathbb{R}}\log\frac{\phi^{\prime\prime}}{e^{-\phi_{0}}}\phi^{\prime\prime}dx.$ Hence,
@@
As shown in proof of [16, Lemma 5, step 2], the last term above equals $\int_{P}v^{\prime\prime}dy$ (as shown by making the change of variables $y=\phi^{\prime}(x)).$ Likewise,
@@ 9.2. The twisted Ding functional of a toric log Fano manifold
(using in the first equality that $\text{(Log)}_{*}\left(|z|^{-2}\frac{i}{2\pi}dz\wedge d\bar{z}\right)=dx$ and [22, Prop 2.9] in the second equality). For example, when $w=0,$ the destabilizing curve for $\mathcal{D}_{\gamma},$ appearing in Cor 4.2 may, if one drops the requirement that it be in $C^{1}(X),$ be expressed explicitly as follows, by taking $v_{0}=0$ in formula 3.37 (which amounts to taking $\phi_{0}(x)\coloneqq|x|)$:
@@
For any metric $g$ on $S^{2}$ compatible with the complex structure on $\mathbb{P}^{1}$
@@
In particular, if $g=g_{S^{2}}$ then $dV_{g}/4\pi$ has unit-mass and thus coincides with $\omega_{\text{FS}}.$ Hence, the defining formula 4.2 for $\Delta_{\text{FS}}$ forces $\Delta_{\text{FS}}=\Delta_{g_{S^{2}}}.$ Next, by formula 3.12,
$E(\mu)=-\frac{1}{2}\int_{X}u_{\mu}\mu,\,\,\text{where}\,dd^{c}(\frac{u_{\mu}}{V})+\omega_{\text{FS}}=\mu,\,\int u_{\mu}\omega_{\text{FS}}=0.$
@@
Finally, in the holomorphic coordinate $z$ on $\mathbb{C}\subset\mathbb{P}^{1}$ we have $dd^{c}(\log(1+|z|^{2})=\omega_{\text{FS}}$ and $dd^{c}(\log|z|^{2})=\delta_{0}.$ Thus
@@
Finally, using stereographic projection gives the explicit formula for $G(\boldsymbol{x},\boldsymbol{y}).$ Indeed, by $SO(3)$-symmetry, the argument of log above equals $(c\left|\boldsymbol{x}-\boldsymbol{y}\right|)^{2}$ for some $c>0$ and using that the diameter of $S^{2}$ equals $2$ forces $c=1/2.$

### v3

@@ 10.1. The free energy of a toric log Fano curve
Denote by $\mathcal{H}_{1,1}(-K_{(X,\Delta)})^{S^{1}}$ the space of all $S^{1}$–invariant psh metrics $\phi$ on $K_{(X,\Delta)}$ such that $\phi(x)$ is in $C^{1,1}(\mathbb{R}).$ Given $\phi_{0}\in\text{PSH(}-K_{(X,\Delta})_{\mathrm{b}}$ which is assumed $S^{1}$–invariant, the corresponding free energy $F_{-\gamma}(\mu)$ satisfies the following formula, when $\mu=dd^{c}\phi/V$ for $\phi\in\mathcal{H}(-K_{(X,\Delta)})_{1,1}^{S^{1}}$:
@@
When $\gamma=1$ it is shown in the proof of [17, Lemma 5, step 1] that
@@
(in our normalization $V\mathcal{M}$ corresponds to the functional $\mathcal{M}$ in [17]). By formula 3.15, this means that
@@
Note that, by formula 10.1, $D(\mu|\mu_{\phi})=\int_{\mathbb{R}}\log\frac{\phi^{\prime\prime}}{e^{-\phi_{0}}}\phi^{\prime\prime}dx.$ Hence,
@@
As shown in proof of [17, Lemma 5, step 2], the last term above equals $\int_{P}v^{\prime\prime}dy$ (as shown by making the change of variables $y=\phi^{\prime}(x)).$ Likewise,
@@ 10.2. The twisted Ding functional of a toric log Fano manifold
(using in the first equality that $\text{(Log)}_{*}\left(|z|^{-2}\frac{i}{2\pi}dz\wedge d\bar{z}\right)=dx$ and [24, Prop 2.9] in the second equality). For example, when $w=0,$ the destabilizing curve for $\mathcal{D}_{\gamma},$ appearing in Cor 4.2 may, if one drops the requirement that it be in $C^{1}(X),$ be expressed explicitly as follows, by taking $v_{0}=0$ in formula 3.37 (which amounts to taking $\phi_{0}(x)\coloneqq|x|)$:
@@
For any Riemannian metric $g$ on $S^{2}$ compatible with the complex structure on $\mathbb{P}^{1}$
@@
In particular, if $g$ is the metric on $S^{2}$ induced by the Euclidean metric on $\mathbb{R}^{3}$ then $dV_{g}/4\pi$ has unit-mass and thus coincides with $\omega_{\text{FS}}.$ Hence, the defining formula 4.2 for $\Delta_{\text{FS}}$ forces $\Delta_{\text{FS}}=\Delta_{g}.$ Next, by formula 3.12,
(10.4) $E(\mu)=-\frac{1}{2}\int_{X}u_{\mu}\mu,\,\,\text{where}\,dd^{c}(\frac{u_{\mu}}{V})=\mu-\omega_{\text{FS}}\,\int u_{\mu}\omega_{\text{FS}}=0.$
@@
Moreover, by formula 3.12 we also have
$E(\mu)=\frac{V}{2}\int_{X}d(\frac{u_{\mu}}{V})\wedge d^{c}(\frac{u_{\mu}}{V})=\frac{V}{2}\left\|\mu-\omega_{\text{FS}}\right\|_{H^{-1}}^{2},$
using the equation for $\frac{u_{\mu}}{V}$ in formula 10.4. Equivalently, integrating by parts,
$E(\mu)=\frac{V}{2}\int_{X}(-dd^{c})(\frac{u_{\mu}}{V})(\frac{u_{\mu}}{V})=\int_{X}(-\Delta_{g})(\frac{u_{\mu}}{V})(\frac{u_{\mu}}{V})dV_{g}=\int_{X}(-\Delta_{g})^{1/2}(\frac{u_{\mu}}{V})(\frac{u_{\mu}}{V})dV_{g}$
which proves the last equality in formula 4.3, using again the equation for $\frac{u_{\mu}}{V}$ in formula 10.4. Finally, to prove the explicit formula for $G(x,y)$ first note that in the holomorphic coordinate $z$ on $\mathbb{C}\subset\mathbb{P}^{1}$ we have $dd^{c}(\log(1+|z|^{2})=\omega_{\text{FS}}$ and $dd^{c}(\log|z|^{2})=\delta_{0}.$ Thus
@@
By stereographic projection this proves the explicit formula for $G(\boldsymbol{x},\boldsymbol{y}).$ Indeed, by $\mathrm{SO}(3)$-symmetry, the argument of log above equals $(c\left|\boldsymbol{x}-\boldsymbol{y}\right|)^{2}$ for some $c>0$ and using that the diameter of $S^{2}$ equals $2$ forces $c=1/2.$